How much potential is required to place a 7×10-3 C charge on a 3000 μF capacitor? 


1
Expert's answer
2021-05-16T18:00:54-0400

"Q = 3000 \times 10^{-6} = 3 \times 10^{-3}\;F \\

C = 7 \times 10^{-3} \;C \\

V = \frac{Q}{C} = \frac{3 \times 10^{-3}}{7 \times 10^{-3}} = 0.428 \;V"